How much heat is required to change 456 g of ice a
Answer #1

Total heat will have three component, first to rise the temperature of ice from -20 to 0, second to melt the ice to water at 0 and then raise the temperature of water from 0 to 25 hence

Q=Q_1+Q_2+Q_3=mS_{ice}(0-(-20))+mL_{fus}+mS_{wat}(25-0)\\ \Rightarrow Q=0.456\times2090\times20+0.456\times33.5\times10^4+0.456\times4186\times25=2.1954\times10^5